# CosmoBench: A Multiscale, Multiview, Multitask Cosmology Benchmark for Geometric Deep Learning

### v1

@@ 5.1 Predicting Cosmological Parameters from Merger Trees
1-Nearest Neighbor Predictor  Our first baseline is a simple nearest-neighbor model. This model considers, for each tree, the empirical distribution of individual features across all nodes and then evaluates a “distance” between trees by computing the Kolmogorov-Smirnov (KS) statistic [69] between empirical distributions. The cosmological parameters $(\Omega_{\rm m},\sigma_{8})$ of each test tree is predicted by the parameters of the closest tree in the training set. More details are given in Sec. C.1. This model does not require any training, however each prediction for a test tree is expensive, and so we restrict the size of the training set to 2,000 trees and the size of the test set to 1,500 trees.
@@
Results and Analysis We report the results on predicting cosmological parameters from a merger tree in the left panel of Tab. 4. We ablate the importance of node features, from individual features including mass $M$, concentration $c$, halo maximum circular velocity $v_{\max}$, and scale factor $a$ (a measure of time), to all of them combined. Notably, the concentration is highly predictive of $\Omega_{\rm m}$, whereas the scale factor $a$ is reasonably predictive of $\sigma_{8}$. Villaescusa-Navarro et al. [12] demonstrated that the internal properties of a single galaxy (mainly stellar mass, stellar metallicity, and maximum circular velocity) are sufficient to predict $\Omega_{\rm m}$ with $\sim 10\%$ precision. Given that present-day galaxy properties encode their formation history, it is reassuring that a single merger tree provides even higher predictive power for $\Omega_{\rm m}$. We find that GNNs offer an edge over DeepSets, showing the advantage of exploiting the tree topology. Smaller models (i.e. hidden size $d=16$) slightly outperform larger on the validation set, likely due to the sparse nature of tree graphs where larger models can overfit.
@@ 5.2 Reconstructing Fine-Scale Merger Trees
Extended Press-Schechter Merger trees can be generated without a cosmological simulation via the Extended Press-Schechter (EPS) formalism [71, 72]. EPS provides a statistical framework for modeling the hierarchical assembly of halos by treating the growth of cosmic structures as a stochastic process, where the overdense regions (halos) effectively undergo a random walk. A merger is considered to occur when this random walk crosses a critical mass threshold, indicating the collapse of a region into a bound structure/halo. EPS starts with a halo at a given time, and probabilistically iterates back in time to construct the merger tree. However, EPS oversimplifies halo growth by assuming that all halos collapse ellipsoidally and that all time steps are uncorrelated. We use the EPS implementation by [73], where the EPS formalism is employed using Monte Carlo sampling, with corrections based on halo mass and shape. To obtain the EPS prediction for a given potentially unresolved merger node, we start the EPS algorithm with the features of the post-merger node (and cosmological parameters associated with the tree), and run the algorithm backwards in time. If the EPS algorithm predicts any split where both halos are above the minimum mass of the two pre-merger nodes, in the time set by the time-steps of the coarsened merger tree, we consider that a prediction of a merger. If not, we take that as the EPS prediction being no merger. Due to the stochastic nature of EPS, we average the predictions over 5 generated trees. We only test it on 20 test trees due to its slow sequential implementation; See Sec. C.2 for more discussion.
@@
Graph Neural Networks We use the same architecture and training set-up as Sec. 5.1, except for changing the batch size as $1$, i.e. taking a gradient step per tree; See Sec. C.2 for more details.
Results and Analysis As shown in Tab. 4 (right), all baselines perform better than random but have vast potential for improvements. We find that the scale factor $a$ is the more discriminative feature for merger node classification, based on which simple $k$-NN based on $1$-hop neighbor information performs competitively to GNN with $4$-hop message-passing. We also note that, unlike EPS, ML methods do not require knowledge of the cosmological parameters, yet still matching or surpassing it.
@@ 6 Conclusion
We presented CosmoBench, a collection of datasets from problems in cosmology designed to benchmark geometric deep learning. These datasets were curated from over 41 million core-hours of simulations and two petabytes of data, and they are currently the largest of their kind. Yet not all of these datasets are large by the standards of ML; many contain only hundreds or thousands of examples and nodes, particularly in CAMELS. We acknowledge this as a limitation and plan to release more data—at smaller scales and larger point clouds—to support future work in more data-rich regimes, better suited to deep architectures. Consequently, there are limits to purely data-driven or “brute-force” approaches, and there are also risks to overfitting with extremely large models. CosmoBench is an invitation to all researchers in ML, working on all types of models, to engage with this data and leverage the geometrical ideas and physical insights to produce enduring solutions in this space.
@@ A.1 Cosmology Glossary
    $N$-body Simulations. A type of cosmological simulation that models the gravitational interaction of a large number of particles (typically representing dark matter). These simulations are computationally intensive and key to studying the growth of structure over time.
@@ A.2 Point Cloud Data Details
For the simulation suites in this work, the particles were initialized using second-order Lagrangian Perturbation theory [^2] https://cosmo.nyu.edu/roman/2LPT/ (2LPT): particles are assigned on a regular cubic periodic grid, and then randomly displaced and assigned peculiar velocities based on the amplitude and shape of the initial power spectrum. This initialization is stochastic—representing the cosmological uncertainty of the initial conditions of the Universe—meaning different point clouds were generated using different initial particle positions and velocities. The values of the cosmological parameters (and the astrophysical ones in the case of CAMELS) are varied over a broad range. The parameter space is sampled using either a Latin-Hypercube or a Sobol Sequence. For Quijote, the parameters varied are $\Omega_{\rm m},\Omega_{\rm b},h,n_{s},\sigma_{8}$. For CAMELS-SAM and CAMELS, the only cosmological parameters varied are $\Omega_{\rm m}$ and $\sigma_{8}$.
@@ A.3 Merger Tree Data Details
Merger trees for the CAMELS-SAM $N$-body simulations were generated in two steps: first, by identifying dark matter halos and subhalos from the particle data using Rockstar; and second by running these halo catalogs through ConsistentTrees [78] to construct the merger trees. The Rockstar algorithm uses an adaptive hierarchical refinement of friends-of-friends groups in seven dimensions (one in time; six in phase-space, positions and velocities). Briefly, friends-of-friends algorithms group two particles if they are within a given distance that scales by the mean particle separation in the initial conditions. The ConsistentTrees algorithm, especially created to pair with Rockstar, was built to explicitly ensure the consistency of halo properties like mass, position, and velocity, across time steps with a novel gravitational method. ConsistentTrees assumes that every halo has at least one progenitor in an earlier time step (logical if assuming large halos build by accretion and merging with other halos), and traces halos backwards in time to check for progenitors in previous timestep catalogs. See Figure 1 of Behroozi et al. [78] for a brief explanation of the algorithm.
@@
To reduce storage requirement and condense information, we trim the trees from CAMELS-SAM by first removing its subtrees with all nodes having mass less than $10^{10}~{}M_{\odot}/h$ (i.e. those with fewer than 100 dark matter particles at all times, a rough guideline for a numerically well-resolved halo). We then prune each path in the resulting tree, where nodes with mass less than $3\times 10^{10}~{}M_{\odot}/h$ are removed, as illustrated in Fig. A.1. This trimming procedure also avoids potential information leakage, as the mass of the smallest halo (which always contains 20 particles) is proportional to $\Omega_{\rm m}$.
@@
In Algorithm 1, we provide the full procedure of constructing the (binary) coarse-grained trees augmented with unresolved merger nodes for classification.
@@
5:     Denote the set of kept nodes $[n^{\prime}]$ and features as $X^{\prime}$.
@@
Figure A.3: Visual comparison of the fine-grained binary tree produced after Pre-Binarization in Algorithm 1 (left), and its corresponding coarse-grained binary tree produced after Post-Binarization in Algorithm 1 (right).
@@ B.1 Predicting Cosmological Parameters from Point Clouds
[…]
@@
For the two-point correlation function cosmology predictions, we construct a simple 4-layer MLP, with tunable hyperparameters. When we pass our two-point correlation function to the MLP, we handle the occasional unphysical (small) negative values at large radii due to shot noise by taking their absolute values. We further scale the correlation features logarithmically to reduce their data range, especially as the values for CAMELS-SAM and CAMELS at lower bins are significantly high. The hidden dimensions of the first and second layers range $[64,128]$, and the third layer from $[16,64]$, with the final layer leading to the prediction of $[\Omega_{\rm m},\sigma_{8}]$. We logarithmically tune the learning rate in the range $[10^{-5},10^{-2}]$, and apply dropout with a rate chosen from the interval $[0.0,0.5]$. The batch sizes range from $\{4,16,64\}$, and the model is trained for a fixed number of 300 epochs. We report the test results from the model selected based on the best validation performance, tuned over 100 trials using the tree-structured Parzen Estimator sampler [79, 80].
@@
We investigate the effect of different 2PCF binning choices: (i) the scale and number of bins for a fixed bin range; (ii) the bin range for fixed scale and number of bins. Ablation (i) is carried out for Quijote using the same bin range between 0.5 and 150 Mpc$/h$, with results shown in Tab. 5. We observe that logarithmic scale is better than linear scale, especially for predicting $\sigma_{8}$ with 25 bins, as it better resolves small distances. Fitting 2PCF-MLP on a logarithmic scale with 25 bins (first row in Tab. 5) produces a test Mean Squared Error (MSE) of $(2.10\pm 0.04)\times 10^{-3},(2.16\pm 0.04)\times 10^{-3}$ for predicting $\Omega_{m},\sigma_{8}$, respectively. Balla et al. [43] reported a test MSE of $(2.03\pm 0.02)\times 10^{-3},(4.66\pm 0.06)\times 10^{-3}$ using the same logrithmic 25 bins and a larger MLP model on a subset of Quijote point clouds (2048/512/512 training/validation/test split). In comparison, using our full Quijote set (19,651/6,551/6,550 training/validation/test split) produces similar performance on predicting $\Omega_{m}$ while notably better performance on predicting $\sigma_{8}$. We skip log feature normalization for Ablation (i) because the Quijote correlation values are within a small range unlike CAMELS-SAM and CAMELS, and we observe no difference by applying log normalization. Ablation (ii) is conducted on all three datasets by comparing 25 logarithmic bins using $R_{\text{min}}=\frac{B}{2000},R_{\text{max}}=\frac{B}{25/3}$ where $B$ is the box size (denoted as “Base” ), with their counterparts that either reduce the minimal bin size four times (denoted as “Base ×$\times\frac{R_{\text{min}}}{4}$”), or quadruple the maximum bin size (denoted as “Base ×$\times 4R_{\text{max}}$”). We report the results in Tab. 6, which shows the effect of the bin range on the performance. We use the choice of “Base ×$\times 4R_{\text{max}}$” for Quijote and CAMELS, and “Base ×$\times\frac{R_{\text{min}}}{4}$” for CAMELS-SAM in the main text Tab. 2.
@@
$\mathbf{R^{2}}$uparrow Bin Choice Quijote
@@
$\mathbf{R^{2}}$uparrow Bin Choice Quijote CAMELS-SAM CAMELS
@@
2PCF Base ×$\times\frac{R_{\text{min}}}{4}$ 0.73 ±$\pm 0.03$ 0.82 ±$\pm 0.02$
Base 0.84 ±$\pm 0.004$ 0.83 ±$\pm 0.004$ 0.64 ±$\pm 0.03$ 0.74 ±$\pm 0.03$ 0.85 ±$\pm 0.02$ 0.30 ±$\pm 0.06$
Base ×$\times 4R_{\text{max}}$ 0.85 ±$\pm 0.004$ 0.84 ±$\pm 0.004$ 0.65 ±$\pm 0.04$ 0.78 ±$\pm 0.03$ 0.84 ±$\pm 0.02$ 0.30 ±$\pm 0.06$
@@
The combinatorial complex is a generalization of graphs, defined as $(S,\chi,\text{rank})$, where $S$ is a set (points in the point cloud), $\chi$ is a skeleton (set of all cells), and $\text{rank}(\cdot)$ maps each cell into its rank in nonnegative integers ($k=0\dots n$). In our experiments, we use tetrahedral cells constructed from a 3-dimensional Delaunay triangulation to identify edge adjacency structure. We also perform ablation study using standard graph with node adjacency (edges) only. See Fig. B.1 for an illustration.
@@
For training the GNNs defined in Equation 2, we perform hyper-parameter search across 100 configurations, including the cutoff radius $R_{c}\in[0.01,0.015,0.02]$ controlling the sparsity of the graph, the types of E(3)-invariant features (Euclidean, Hausdorff, None), the number of GNN layers $L\in\{1,\ldots,6\}$, and the GNN hidden dimensions from $\{32,64,128,256\}$. We explore different intra- and inter-aggregation functions (i.e., sum, max, min, and all three combined) and nonlinear activation functions (tanh or ReLU). For datasets CAMELS-SAM and CAMELS, we limit the batch size to smaller values ($\{1,2,4,8\}$), while for the Quijote we include batch sizes up to 16. We also search over the learning rate $[10^{-5},10^{-3}]$, weight decay $[10^{-7},10^{-5}]$, neighborhood dropout probability $[0,0.2]$, and the number of epochs for cosine annealing ($T_{\text{max}}\in\{10,100\}$). Each trial consists of 300 epochs, and we report our test results based on the best validation results. We conduct 100 trials for both CAMELS-SAM and CAMELS, whereas for Quijote, we carry out 25 trials due to computational limitations.
@@
Table 7: GNN Ablation on Cosmological Parameter Regression ($R^{2}\pm$ 1std). Time unit: 1-GPU time.
$\mathbf{R^{2}}$uparrow Quijote CAMELS-SAM CAMELS
GNN 0.80 ±$\pm 0.004$ 0.77±$\pm 0.005$ 671K 1 day 0.75 ±$\pm 0.03$ 0.83 ±$\pm 0.02$ 1003K 3 hr 0.78 ±$\pm 0.03$ 0.24 ±$\pm 0.06$ 1166K 2 hr
GNN (w/o edgeMP) 0.80 ±$\pm 0.004$ 0.79±$\pm 0.005$ 128K 1 day 0.72 ±$\pm 0.03$ 0.84 ±$\pm 0.02$ 506K 3 hr 0.80 ±$\pm 0.02$ 0.27 ±$\pm 0.06$ 384K 2 hr
We perform an ablation study to see the impact of edge-edge message-passing defined on the higher-order graphs. We adopt the GNN (w/o edgeMP) model by disabling message-passing between edges defined on the parent tetrahedron (see Fig. B.1 as a comparison). We observe marginal differences in the performance metrics mostly within the 1 std, as shown in Tab. 7 and Fig. B.2. This may arise from multiple factors, including optimization challenges introduced by increased model complexity, or the possibility that the constructed higher-order graphs do not offer structural advantages over simpler ones. As a proof of concept, our higher-order graphs operate on tetrahedral cells identified from the Delaunay triangulation. We note that such tetrahedral cells are still locally confined, not modeling much larger scales compared to edges in standard graphs. Incorporating hyper-edges or clusters beyond edges or tetrahedral cells may further improve the performance on cosmological parameter prediction as shown recently in [63]. Our code base [^3] https://github.com/Byeol-Haneul/CosmoTopo/tree/benchmark allows exploration to even higher-orders involving a large number of galaxies and halos, at diverse scales. In future work, we aim to improve the construction of combinatorial complexes adapted to these cosmological point clouds for better performance.
@@
Our results reported in Tab. 2 (with further details in Fig. B.2) show that ML baselines are comparable to 2PCF baseline for Quijote and CAMELS, while being strictly better for CAMELS-SAM. A promising next step for a more comprehensive evaluation is to test these baselines on observational data which include experimental noise.

### v2

@@ 5.1 Predicting Cosmological Parameters from Merger Trees
1-Nearest Neighbor Predictor Our first baseline is a simple nearest-neighbor model. This model considers, for each tree, the empirical distribution of individual features across all nodes and then evaluates a “distance” between trees by computing the Kolmogorov-Smirnov (KS) statistic [69] between empirical distributions. The cosmological parameters $(\Omega_{\rm m},\sigma_{8})$ of each test tree is predicted by the parameters of the closest tree in the training set. More details are given in Sec.˜C.1. This model does not require any training, however each prediction for a test tree is expensive, and so we restrict the size of the training set to 2,000 trees and the size of the test set to 1,500 trees.
@@
Results and Analysis We report the results on predicting cosmological parameters from a merger tree in the left panel of Tab.˜5. We ablate the importance of node features, from individual features including mass $M$, concentration $c$, halo maximum circular velocity $v_{\max}$, and scale factor $a$ (a measure of time), to all of them combined. Notably, the concentration is highly predictive of $\Omega_{\rm m}$, whereas the scale factor $a$ is reasonably predictive of $\sigma_{8}$. Villaescusa-Navarro et al. [12] demonstrated that the internal properties of a single galaxy (mainly stellar mass, stellar metallicity, and maximum circular velocity) are sufficient to predict $\Omega_{\rm m}$ with $\sim 10\%$ precision. Given that present-day galaxy properties encode their formation history, it is reassuring that a single merger tree provides even higher predictive power for $\Omega_{\rm m}$. We find that GNNs offer an edge over DeepSets, showing the advantage of exploiting the tree topology. Smaller models (i.e. hidden size $d=16$) slightly outperform larger on the validation set, likely due to the sparse nature of tree graphs where larger models can overfit.
@@ 5.2 Reconstructing Fine-Scale Merger Trees
Extended Press-Schechter Merger trees can be generated without a cosmological simulation via the Extended Press-Schechter (EPS) formalism [71, 72]. EPS provides a statistical framework for modeling the hierarchical assembly of halos by treating the growth of cosmic structures as a stochastic process, where the overdense regions (halos) effectively undergo a random walk. A merger is considered to occur when this random walk crosses a critical mass threshold, indicating the collapse of a region into a bound structure/halo. EPS starts with a halo at a given time, and probabilistically iterates back in time to construct the merger tree. However, EPS oversimplifies halo growth by assuming that all halos collapse ellipsoidally and that all time steps are uncorrelated. We use the EPS implementation by [73], where the EPS formalism is employed using Monte Carlo sampling, with corrections based on halo mass and shape. To obtain the EPS prediction for a given potentially unresolved merger node, we start the EPS algorithm with the features of the post-merger node (and cosmological parameters associated with the tree), and run the algorithm backwards in time. If the EPS algorithm predicts any split where both halos are above the minimum mass of the two pre-merger nodes, in the time set by the time-steps of the coarsened merger tree, we consider that a prediction of a merger. If not, we take that as the EPS prediction being no merger. Due to the stochastic nature of EPS, we average the predictions over 5 generated trees. We only test it on 20 test trees due to its slow sequential implementation; See Sec.˜C.2 for more discussion.
@@
Graph Neural Networks We use the same architecture and training set-up as Sec.˜5.1, except for changing the batch size as $1$, i.e. taking a gradient step per tree; See Sec.˜C.2 for more details.
Results and Analysis As shown in Tab.˜5 (right), all baselines perform better than random but have vast potential for improvements. We find that the scale factor $a$ is the more discriminative feature for merger node classification, based on which simple $k$-NN based on $1$-hop neighbor information performs competitively to GNN with $4$-hop message-passing. We also note that, unlike EPS, ML methods do not require knowledge of the cosmological parameters, yet still matching or surpassing it.
@@ 6 Conclusion
We present CosmoBench, a collection of datasets from problems in cosmology designed to benchmark geometric deep learning. These datasets are curated from over 41 million core-hours of simulations and two petabytes of data, and they are currently the largest of their kind. Yet not all of these datasets are large by the standards of ML; many contain only hundreds or thousands of examples and nodes, particularly in CAMELS. Consequently, there are limits to purely data-driven or “brute-force” approaches, and there are also risks to overfitting with extremely large models. CosmoBench is an invitation to all researchers in ML, working on all types of models, to engage with this data and leverage the geometrical ideas and physical insights to produce enduring solutions in this space.
Looking ahead, we plan to expand CosmoBench with more data—at smaller scales and larger point clouds—to support development of deep architectures and foundation models in more data-rich regimes. Another promising direction is to utilize ML emulators to produce simulation data more efficiently, complementing traditional computationally-intensive simulation pipelines. To bridge the gap between simulations and observational data, we have taken the first step by including mock observations in redshift space for the velocity prediction task, and aim to further develop our benchmark with more survey realism and observational challenges.
Acknowledgments and Disclosure of Funding
The authors thank David Hogg (Flatiron), Soledad Villar (JHU), Jahmour Givansand (Flatiron), Natalí de Santi (UCB), Bruno Régaldo-Saint Blancard (Flatiron), Ruben Ohana (NVIDIA), Chirag Modi (Flatiron), and David Spergel (Flatiron) for valuable discussions, as well as Maya Bechler-Speicher (Meta), Michael Galkin (Google), and Christopher Morris (RWTH Aachen) for motivating this work. The authors also thank the anonymous NeurIPS reviewers and area chair for their constructive feedback. RS acknowledges financial support from STFC Grant No. ST/X508664/1, the Snell Exhibition of Balliol College, Oxford, and the CCA Pre-doctoral Program.
@@ A.1 Cosmology Glossary
    NN-body Simulations. A type of cosmological simulation that models the gravitational interaction of a large number of particles (typically representing dark matter). These simulations are computationally intensive and key to studying the growth of structure over time.
@@ A.2 Point Cloud Data Details
For the simulation suites in this work, the particles are initialized using second-order Lagrangian Perturbation theory [^3] https://cosmo.nyu.edu/roman/2LPT/ (2LPT): particles are assigned on a regular cubic periodic grid, and then randomly displaced and assigned peculiar velocities based on the amplitude and shape of the initial power spectrum. This initialization is stochastic—representing the cosmological uncertainty of the initial conditions of the Universe—meaning different point clouds were generated using different initial particle positions and velocities. The values of the cosmological parameters (and the astrophysical ones in the case of CAMELS) are varied over a broad range. The parameter space is sampled using either a Latin-Hypercube or a Sobol Sequence. For Quijote, the parameters varied are $\Omega_{\rm m},\Omega_{\rm b},h,n_{s},\sigma_{8}$. For CAMELS-SAM and CAMELS, the only cosmological parameters varied are $\Omega_{\rm m}$ and $\sigma_{8}$.
@@ A.3 Merger Tree Data Details
Merger trees for the CAMELS-SAM $N$-body simulations are generated in two steps: first, by identifying dark matter halos and subhalos from the particle data using Rockstar; and second by running these halo catalogs through ConsistentTrees [78] to construct the merger trees. The Rockstar algorithm uses an adaptive hierarchical refinement of friends-of-friends groups in seven dimensions (one in time; six in phase-space, positions and velocities). Briefly, friends-of-friends algorithms group two particles if they are within a given distance that scales by the mean particle separation in the initial conditions. The ConsistentTrees algorithm, especially created to pair with Rockstar, was built to explicitly ensure the consistency of halo properties like mass, position, and velocity, across time steps with a novel gravitational method. ConsistentTrees assumes that every halo has at least one progenitor in an earlier time step (logical if assuming large halos build by accretion and merging with other halos), and traces halos backwards in time to check for progenitors in previous timestep catalogs. See Figure 1 of Behroozi et al. [78] for a brief explanation of the algorithm.
@@
To reduce storage requirement, condense information, and prevent information leakage, we trim the trees from CAMELS-SAM by first removing its subtrees with all nodes having mass less than $10^{10}~M_{\odot}/h$ (i.e. those with fewer than 100 dark matter particles at all times, a rough guideline for a numerically well-resolved halo). We then prune each path in the resulting tree, where nodes with mass less than $3\times 10^{10}~M_{\odot}/h$ are removed, as illustrated in Fig.˜A.1. This trimming procedure also avoids potential information leakage, as the mass of the smallest halo (which always contains 20 particles) is proportional to $\Omega_{\rm m}$.
@@
In Algorithm˜1, we provide the full procedure of constructing the (binary) coarse-grained trees augmented with unresolved merger nodes for classification.
@@
5:   Denote the set of kept nodes $n^{\prime}]$ and features as $X^{\prime}$.
@@
Figure A.3: Visual comparison of the fine-grained binary tree produced after Pre-Binarization in Algorithm˜1 (left), and its corresponding coarse-grained binary tree produced after Post-Binarization in Algorithm˜1 (right).
@@ B.1 Predicting Cosmological Parameters from Point Clouds
For all datasets, we adopt logarithmic radial binning. For the Quijote halos, simulated in a 1000 Mpc/h/h box, we use 25 bins spaced logarithmically between 0.5 and 120 Mpc/h/h. For the CAMELS-SAM galaxy catalog, extracted from a 100 Mpc/h/h box, we use 25 bins spanning 0.0125 to 12 Mpc/h/h. For the higher-resolution CAMELS sample, which resides in a 25 Mpc/h/h box, we use 25 bins between 0.0125 and 3 Mpc/h/h. To compute $\mathrm{DR},\mathrm{RR}$ in Equation Equation˜9, we use a random point cloud with 100 times more points than the data. The $\{\xi_{i}\}$ then become the input features for the MLP, as we discuss next.
@@
For the two-point correlation function cosmology predictions, we construct a simple 4-layer MLP, with tunable hyperparameters. When we pass our two-point correlation function to the MLP, we handle the occasional unphysical (small) negative values at large radii due to shot noise by taking their absolute values. We further scale the correlation features logarithmically to reduce their data range, especially as the values for CAMELS-SAM and CAMELS at lower bins are significantly high. The hidden dimensions of the first and second layers range $[64,128]$, and the third layer from $[16,64]$, with the final layer leading to the prediction of $\Omega_{\rm m},\sigma_{8}]$. We logarithmically tune the learning rate in the range $10^{-5},10^{-2}]$, and apply dropout with a rate chosen from the interval $[0.0,0.5]$. The batch sizes range from $\{4,16,64\}$, and the model is trained for a fixed number of 300 epochs. We report the test results from the model selected based on the best validation performance, tuned over 100 trials using the tree-structured Parzen Estimator sampler [79, 80].
@@
We investigate the effect of different 2PCF binning choices: (i) the scale and number of bins for a fixed bin range; (ii) the bin range for fixed scale and number of bins. Ablation (i) is carried out for Quijote using the same bin range between 0.5 and 150 Mpc/h/h, with results shown in Tab.˜6. We observe that logarithmic scale is better than linear scale, especially for predicting $\sigma_{8}$ with 25 bins, as it better resolves small distances. Fitting 2PCF-MLP on a logarithmic scale with 25 bins (first row in Tab.˜6) produces a test Mean Squared Error (MSE) of $(2.10\pm 0.04)\times 10^{-3},(2.16\pm 0.04)\times 10^{-3}$ for predicting $\Omega_{\rm m},\sigma_{8}$, respectively. Balla et al. [43] reported a test MSE of $(2.03\pm 0.02)\times 10^{-3},(4.66\pm 0.06)\times 10^{-3}$ using the same logrithmic 25 bins and a larger MLP model on a subset of Quijote point clouds (2048/512/512 training/validation/test split). In comparison, using our full Quijote set (19,651/6,551/6,550 training/validation/test split) produces similar performance on predicting $\Omega_{\rm m}$ while notably better performance on predicting $\sigma_{8}$. We skip log feature normalization for Ablation (i) because the Quijote correlation values are within a small range unlike CAMELS-SAM and CAMELS, and we observe no difference by applying log normalization. Ablation (ii) is conducted on all three datasets by comparing 25 logarithmic bins using $R_{\text{min}}=\frac{B}{2000},R_{\text{max}}=\frac{B}{25/3}$ where $B$ is the box size (denoted as “Base” ), with their counterparts that either reduce the minimal bin size four times (denoted as “Base $\times\frac{R_{\text{min}}}{4}$”), or quadruple the maximum bin size (denoted as “Base $\times 4R_{\text{max}}$”). We report the results in Tab.˜7, which shows the effect of the bin range on the performance. We use the choice of “Base $\times 4R_{\text{max}}$” for Quijote and CAMELS, and “Base $\times\frac{R_{\text{min}}}{4}$” for CAMELS-SAM in the main text Tab.˜2.
@@
$\mathbf{R^{2}}$ $\uparrow$ Bin Choice Quijote
@@
$\mathbf{R^{2}}$ $\uparrow$ Bin Choice Quijote CAMELS-SAM CAMELS
@@
2PCF Base $\times\frac{R_{\text{min}}}{4}$ 0.73 $\pm 0.03$ 0.82 $\pm 0.02$
Base 0.84 $\pm 0.004$ 0.83 $\pm 0.004$ 0.64 $\pm 0.03$ 0.74 $\pm 0.03$ 0.85 $\pm 0.02$ 0.30 $\pm 0.06$
Base $\times 4R_{\text{max}}$ 0.85 $\pm 0.004$ 0.84 $\pm 0.004$ 0.65 $\pm 0.04$ 0.78 $\pm 0.03$ 0.84 $\pm 0.02$ 0.30 $\pm 0.06$
LLS Ablation
We investigate the effect of the cutoff radius in LLS, which are selected using the validation set to greedily find the 12 best cutoff thresholds given a candidate threshold set, and chosen separately for each target parameter $(\Omega_{\rm m},\sigma_{8})$. We now provide another LLS baseline on a “naive” set of cutoff features, which are chosen from the candidate set with equi-spaced elements, and applied to both parameters. As shown in Table 8, this naive choice of cutoff features (first row) yields a poorer fit from the linear model than the optimized features used in the main paper (second row) and performs worse than GNNs (third row) in CAMELS-SAM for both parameters and CAMELS for $\Omega_{\rm m}$.
Table 8: The Effect of LLS Features for Cosmological Parameter Regression
$\mathbf{R^{2}}$ $\uparrow$ Quijote CAMELS-SAM CAMELS
$\Omega_{\rm m}$ $\sigma_{8}$ $\Omega_{\rm m}$ $\sigma_{8}$ $\Omega_{\rm m}$ $\sigma_{8}$
LLS (naive) 0.83 $\pm$0.004 0.80 $\pm$0.004 0.68 $\pm$0.04 0.73 $\pm$0.03 0.74 $\pm$0.03 0.26 $\pm$0.06
LLS 0.83 $\pm$0.004 0.80 $\pm$0.004 0.77 $\pm$0.03 0.82 $\pm$0.02 0.78 $\pm$0.03 0.28 $\pm$0.06
GNN 0.80 $\pm$0.004 0.77 $\pm$0.005 0.75 $\pm$0.03 0.83 $\pm$0.02 0.78 $\pm$0.03 0.24 $\pm$0.06
@@
The combinatorial complex is a generalization of graphs, defined as $(S,\chi,\text{rank})$, where $S$ is a set (points in the point cloud), $\chi$ is a skeleton (set of all cells), and $\text{rank}(\cdot)$ maps each cell into its rank in nonnegative integers ($k=0\dots n$). In our experiments, we use tetrahedral cells constructed from a 3-dimensional Delaunay triangulation to identify edge adjacency structure. We also perform ablation study using standard graph with node adjacency (edges) only. See Fig.˜B.1 for an illustration.
@@
For training the GNNs defined in Equation˜2, we perform hyper-parameter search across 100 configurations, including the cutoff radius $R_{c}\in[0.01,0.015,0.02]$ controlling the sparsity of the graph, the types of E(3)-invariant features (Euclidean, Hausdorff, None), the number of GNN layers $L\in\{1,\ldots,6\}$, and the GNN hidden dimensions from $\{32,64,128,256\}$. We explore different intra- and inter-aggregation functions (i.e., sum, max, min, and all three combined) and nonlinear activation functions (tanh or ReLU). For datasets CAMELS-SAM and CAMELS, we limit the batch size to smaller values ($\{1,2,4,8\}$), while for the Quijote we include batch sizes up to 16. We also search over the learning rate $10^{-5},10^{-3}]$, weight decay $10^{-7},10^{-5}]$, neighborhood dropout probability $[0,0.2]$, and the number of epochs for cosine annealing ($T_{\text{max}}\in\{10,100\}$). Each trial consists of 300 epochs, and we report our test results based on the best validation results. We conduct 100 trials for both CAMELS-SAM and CAMELS, whereas for Quijote, we carry out 25 trials due to computational limitations.
@@
We perform an ablation study to see the impact of edge-edge message-passing defined on the higher-order graphs. We adopt the GNN (w/o edgeMP) model by disabling message-passing between edges defined on the parent tetrahedron (see Fig.˜B.1 as a comparison). We observe marginal differences in the performance metrics mostly within the 1 std, as shown in Tab.˜2 and Fig.˜B.2. This may arise from multiple factors, including optimization challenges introduced by increased model complexity, or the possibility that the constructed higher-order graphs do not offer structural advantages over simpler ones. As a proof of concept, our higher-order graphs operate on tetrahedral cells identified from the Delaunay triangulation. We note that such tetrahedral cells are still locally confined, not modeling much larger scales compared to edges in standard graphs. Incorporating hyper-edges or clusters beyond edges or tetrahedral cells may further improve the performance on cosmological parameter prediction as shown recently in [63]. Our code base [^4] https://github.com/Byeol-Haneul/CosmoTopo/tree/benchmark allows exploration to even higher-orders involving a large number of galaxies and halos, at diverse scales. In future work, we aim to improve the construction of combinatorial complexes adapted to these cosmological point clouds for better performance.
@@
Our results reported in Tab.˜2 (with further details in Fig.˜B.2) show that ML baselines are comparable to 2PCF baseline for Quijote and CAMELS, while being strictly better for CAMELS-SAM. A promising next step for a more comprehensive evaluation is to test these baselines on observational data which include experimental noise.
Predicting Additional Parameters.
For our baseline tasks, we focus on the two cosmological parameters $\Omega_{\rm m},\sigma_{8}$ due to their significance in astrophysics (they are among the least well-constrained cosmological parameters from the observational data), as well as the feasibility of predicting them from the positions of halos and galaxies (other parameters such as $\Omega_{\rm b},n_{s},h$ are more difficult to measure from positions alone and typically require information such as the cosmic microwave background or supernovae). Nonetheless, we are interested to see how well ML can predict additional parameters and bridge the gap to real-world inference problems. A shown in Tab.˜9, these baselines perform significantly worse on predicting $\Omega_{\rm b},n_{s},h$ compared to $\Omega_{\rm m},\sigma_{8}$, as expected.
Table 9: Cosmological Parameter Regression using Quijote: $\Omega_{\rm b}$, $n_{s}$, and $h$.
R2 $\uparrow$ $\Omega_{\rm b}$ $n_{s}$ $h$
[…]